OpenStudy (anonymous):

find the geometric power series for the function \[\frac{ 1 }{ 2+x }\] centered at 0

OpenStudy (anonymous):

as you may recall: \[\sum_{0}^{\infty} x^{n} = \frac{ 1 }{ 1 - x }\] we want to put the function of x in the form of the formula so we can simply plug it in the series.

OpenStudy (anonymous):

let me know if i should do the next step

OpenStudy (anonymous):

im thinking the answer is \[\sum_{n=0}^{\infty} \frac{ -x ^{n} }{ 2^{n+1} }\] ?

OpenStudy (anonymous):

it is. way to simplify. lol

OpenStudy (anonymous):

oh, thank you!